# VFScale: Intrinsic Reasoning through Verifier-Free Test-time Scalable Diffusion Model

### v3

@@ Appendix F Broader Impacts
This paper aims to advance machine learning—particularly diffusion-based generative models. While our improvements promise higher-quality AI-generated content, they also carry potential societal risks. Accordingly, we must remain vigilant to unintended negative outcomes and guard against any unethical or illicit applications of this technology.
Appendix G Limitations and future work
@@
Our inference framework primarily relies on MCTS, which presents two key limitations: (1) limited compatibility with parallel computing, and (2) challenges in effectively evaluating node quality during the early stages of denoising. Future work could explore integrating alternative search strategies, such as those proposed by [28]. Additionally, to enhance performance-energy consistency, we introduce linear-regression negative contrastive learning, which enforces a linear relationship between energy and the distance to real samples. Further investigation is needed to assess the broader implications of this constraint and explore alternative regularization approaches. Lastly, while our current implementation utilizes Gaussian noise for branching, other diffusion-based branching mechanisms remain an open area for exploration.
Appendix H Visualization of results
H.1 Visualization of Maze experiments
@@
H.2 Visualization of Sudoku experiments
@@
This section presents visualizations of the training and test Sudoku data in Fig. 12, and representative samples generated by different methods in Fig. 13. In the visuals, black numbers denote the condition, green numbers represent correct predictions, and red numbers represent wrong predictions. All visualizations are derived from a few representative samples. The comparison between the training and test sets clearly indicates that the tasks in the test set are significantly more difficult than those in the training set. When comparing the samples generated by different methods, it is evident that the originally trained model, regardless of the inference strategy, consistently underperforms compared to VFScale.

### v4

@@ Appendix G Broader Impacts
This paper aims to advance machine learning, particularly diffusion-based generative models. While our improvements promise higher-quality AI-generated content, they also carry potential societal risks. Accordingly, we must remain vigilant to unintended negative outcomes and guard against any unethical or illicit applications of this technology.
Appendix H The Use of LLMs
In the preparation of this manuscript, Large Language Models (LLMs) were used as an assistive tool to improve the quality of the work. The specific uses include:
  • Writing and Language Refinement: LLMs were utilized to paraphrase sentences for clarity, correct grammatical errors, and refine the overall language to meet high academic standards.
  • Coding and Engineering Assistance: LLMs were employed to generate boilerplate code, assist in debugging, and accelerate the implementation of standard algorithms, which supported the engineering of our experimental framework.
The authors reviewed, edited, and validated all LLM-generated content (both text and code) to ensure its technical accuracy and originality. The authors take full responsibility for all content presented in this paper.
Appendix I Limitations and future work
@@
Our inference framework primarily relies on MCTS, which presents two key limitations: (1) limited compatibility with parallel computing, and (2) challenges in effectively evaluating node quality during the early stages of denoising. Future work could explore integrating alternative search strategies, such as those proposed by Wu et al. (2024). Additionally, to enhance performance-energy consistency, we introduce linear-regression negative contrastive learning, which enforces a linear relationship between energy and the distance to real samples. Further investigation is needed to assess the broader implications of this constraint and explore alternative regularization approaches. Lastly, while our current implementation utilizes Gaussian noise for branching, other diffusion-based branching mechanisms remain an open area for exploration.
Appendix J Visualization of results
J.1 Visualization of Maze experiments
@@
J.2 Visualization of Sudoku experiments
@@
This section presents visualizations of the training and test Sudoku data in Fig. 12, and representative samples generated by different methods in Fig. 13. In the visuals, black numbers denote the condition, green numbers represent correct predictions, and red numbers represent wrong predictions. All visualizations are derived from a few representative samples. The comparison between the training and test sets clearly indicates that the tasks in the test set are significantly more difficult than those in the training set. When comparing the samples generated by different methods, it is evident that the originally trained model, regardless of the inference strategy, consistently underperforms VFScale.
Appendix K Ethics Statement
All authors have read and adhered to the ICLR Code of Ethics. This research focuses on the foundational capabilities of generative models for algorithmic reasoning tasks. The datasets used are synthetically generated and do not contain personally identifiable information or sensitive data. Our work does not involve human subjects. While our primary contribution is algorithmic, we acknowledge that advancements in generative models can have broader societal impacts. We believe the potential for misuse of this specific technology is low, but we encourage the responsible development and application of all AI systems.
Appendix L Reproducibility statement
To ensure the reproducibility of our research, we have made the complete source code, datasets, and pre-trained model checkpoints publicly available at https://github.com/AI4Science-WestlakeU/VFScale. The repository contains the full implementation of our proposed VFScale framework, including all training objectives (e.g., MRNCL) and the hMCTS inference algorithm. Furthermore, comprehensive details regarding our experimental setup, model architectures, hyperparameter settings, and computational resources are provided in the Appendix.